OpenStudy (anonymous):

Solve for x. -3 + 8x - 5 = -8

OpenStudy (anonymous):

8x=0 X=0

OpenStudy (anonymous):

-3 + 8x - 5 = -8 +3-3+8x-5+5=-8+5+3 8x=0 x=0